OpenStudy (anonymous):

use the sq root proprty to solve (x+5)^2=10 is the answer 5+ or - sq root 10

OpenStudy (phi):

almost \[ (x+5)^2 =10\] take the square root of both sides \[\sqrt{(x+5)^2}= \sqrt{10} \] of course the square root undoes the square, so this is \[ x+5= \sqrt{10} \] add -5 to both sides \[ x+5-5= -5+\sqrt{10} \] simplify 5-5=0 and x+0 is just x , so \[ x= -5 + \sqrt{10} \] next we remember that -sqrt(10) is also the square root of 10, so the other answer is \[ x= -5 - \sqrt{10} \] we can write this in short hand using ± : \[ x= -5 ± \sqrt{10} \]

OpenStudy (anonymous):

ok now i understand thank you so mcuh